OpenStudy (anonymous):

Find the remainder when 11!+9! (999*8!) is divided by 18*7!-14*6! (2*8!) ?

ganeshie8 (ganeshie8):

usually the answer is 0 for these type of questions

ganeshie8 (ganeshie8):

First, notice that the given expression is divisible by 8!

ganeshie8 (ganeshie8):

11!+9! (999*8!) = 8!(9*10*11 + 9!*999)

OpenStudy (anonymous):

I the solution it's given the ans to be 1*8! = 8!

ganeshie8 (ganeshie8):

we can factor out a "2" too : 11!+9! (999*8!) = 8!(9*10*11 + 9!*999) = 2*8!(9*5*11 + (9!/2)*999)

OpenStudy (anonymous):

in the*

OpenStudy (anonymous):

@ganeshie8 Yes i under stand that, 8! and 8! in the num and Denominator cancells out. and remaining is 999/2

OpenStudy (anonymous):

But in the solution it's given 1*8! and the book i am referring is never wrong, its a standard book

ganeshie8 (ganeshie8):

Looks I have misinterpreted the problem, one sec..

OpenStudy (anonymous):

@ganeshie8 Let me post a screenshot

ganeshie8 (ganeshie8):

that would help, thank you :)

OpenStudy (anonymous):

@ganeshie8 Here it is

OpenStudy (anonymous):

@ganeshie8 Any Clue's ?

ganeshie8 (ganeshie8):

Ahh this should be easy..

ganeshie8 (ganeshie8):

11!+9! 8!(999) 8!(998+1) 8!*998 + 8!*1 8!*2*499 + 8!

ganeshie8 (ganeshie8):

Notice that 8!*2*499 is divisible by 8!*2 therefore the remainder is just 8!

OpenStudy (anonymous):

@ganeshie8 (8!*2*499+8!) / (2*8!)

OpenStudy (anonymous):

8!*2*499 is divisible by 8!*2 , buy here 8! and 8! are getting cancelled and also the 2, left one is 499

OpenStudy (anonymous):

@ganeshie8 Please make me understand.

ganeshie8 (ganeshie8):

Hey! tag me once you're online... I'd like to help you understand how these work :)